34th HiLumi WP2 Task Leader Meeting
Minutes of the 34th WP2 Task Leader Meeting held on 17/09/2014
Present: G. Arduini, R. De Maria, M. Giovannozzi, S. Fartoukh, E. Métral, R. Tomas, A. Valishev.
Minutes, Follow-up of Actions, General Information (Gianluigi)
The last minutes have been approved with comments from Massimo.
A meeting on magnet powering with Ezio, Amalia, Jean-Paul will be organized on the 26th of September. Gianluigi thanked Massimo and Riccardo for organizing that.
Follow-up and new actions:
- Gianluigi found 2.4 W/m/aperture as cooling capacity for the beam screen in the arcs in a presentation given by S. Claudet (compared to 1W/m/aperture for Q7 from impedance assuming the most pessimistic case – 50 ns beam – and the highest operational temperature – 70 K, see table attached https://indico.cern.ch/event/323863/contribution/3/material/slides/1.pdf), but this is the limit due to the dimensions of the beam screen capillaries and might not be achievable with the existing cryo-plants. To be included: e-cloud and synchrotron radiation. A talk at the annual meeting should review the total expected heat loads in the beam screen for HL-LHC. Gianluigi asked Elias to verify with Serge to see the real limitations for HL-LHC as part of the Task 2.7. Action: Elias.
After the meeting Elias checked with S. Claudet that:
- Local limit (due to the dimensions of the beam screen capillaries): 2.4 W/m/aperture => 4.8 W/m.
- Global limit for the cryo-plant: 3.8 W/m => This is for instance the limit we already used in our talk at hilumi meeting at Darebury in 2013 (slide 15 of https://indico.cern.ch/event/257368/session/11/contribution/33/material/slides/1.pdf).
- S. SergeC is trying to restart at the beginning of 2015 a “Heat load working group” to see how we share all the head load contributions (impedance, e-cloud, synchrotron radiation, and eddy currents) => Would be perfect for the interface with Task 2.7.
- Gianluigi reminded the pending action to compare the DA aperture without and with beam-beam for HL-LHCV1.0 in the with/without magnetic field errors. Massimo mentioned that the simulations for latest version of the error tables are ongoing at SLAC. Sasha and Tatiana should make sure that they use the same error tables. Action: Sasha and Tatiana to check with Massimo. Gianluigi noted that the impact of the individual multipoles should be studied also with beam-beam if an impact on the dynamic aperture is observed. This will be presented at the KEK workshop but it is important to have a status at one of the WP2 Task Leaders meetings.
- For 6D lens modification in SixTrack a status report is expected for end of October. Action: Tatiana.
- For the annual meeting the program is being prepared. The plenary talks foreseen are: layout and integration, optics and operation modes, QXF cable review, collimation status, machine protection (R2EE and availability workshop), LIU progress, outcome on detector, status on PDR, alternative scenario talk, evolution on the collaboration governance, evolution on the LARP program. Friday afternoon there will be a task leader meeting dedicated to the step required for the Cost and Schedule Review and for the TDR.
Down selection criteria and MDs prior to LS 3 (Gianluigi)
A list of MDs has been identified for validating the nominal scenario:
- 25ns operation: study the scrubbing process in the LHC, evaluation of the performance of a-C coating at cryogenic temperature with the COLDEX experiment in the SPS. Stephane asked about the size of the COLDEX beam screen and the type of measurements that will be done. The diameter of the beam screen is 70 mm and the instrumentation and experiments are listed in the slides.
- Crab cavities: validating in the SPS cavity control, noise issues, stealth mode. It is foreseen to use the crab cavities in counter-phasing at low voltage at injection energy. Stephane asked what expected orbit deviation is in the region of the cavities. Riccardo replied that this should be of the order of 5 mm (for two distant cavities with kick phase 0.5 MV(e) / 450GeV 6.6 m results in 7.3 um or for closer cavities 0.5 MV(e) / 450GeV 1.1 m results in 1.2 um).
- b* levelling is the nominal scenario for luminosity levelling and the principle has been demonstrated in MDs but orbit/separation reproducibility need to be assessed as well as its robustness against instabilities. Stephane pointed out that levelling with parallel separation should not be excluded and systematic tests to study beam stability as function of separation with both octupole polarities should be conducted. Gianluigi fully agreed and it will be added to the list of experiments to be performed.
- Beam-Beam Long Range Compensator could be used to reduce the crossing angle to mitigate possible performance limitations of the crab cavities. Stephane noted that the long range compensator could allow avoiding the alternate crossing in IP1/5. In particular vertical-vertical crossing would be interesting from the energy deposition point of view.
Stephane and Rogelio asked whether the scheduling and prioritization of the MD should be discussed. Gianluigi replied that this should be done in a second stage the aim of the presentation being to present the criteria and studies required for down-selection.
The levelling methods triggered a discussion on the simulation of beam-beam effect and the impact of the field errors on the DA even in the presence of beam-beam. Stephane pointed out that this is worrying, Gianluigi agreed and noted that an action on the evaluation of the impact of field quality in the presence of beam-beam was ongoing (see above).
Updates from Task Leaders (Task leaders)
- Task 2.2: Barbara progressed on the optics with an additional Q7 to enhance the b function at the crab cavities.
- Task 2.3: Additional simulations concerning the D2 field quality have been presented at a WP3 meeting. Ezio should circulate the corresponding field quality tables for tracking checks.
- Task 2.4: Studies on the inter-play between beam-beam and impedance are ongoing. A summary should be presented at the WP2 Task Leader meeting.
- Task 2.5: Work is ongoing to implement noise and IBS in Lifetrack, discussion with Herman on electron lens are ongoing.
Next meeting will take place on Friday 3rd October 2014 in collaboration with WP2/WP5/WP10. For Friday 10th October 2014, Tatiana is expected to present PACMAN effects and Ezio is going to give an update on the magnet field quality.
Reported by Gianluigi and Riccardo.